solve the system of equations by substitution
y=3x+7
y=x-9

Answer by jim_thompson5910(35256)


y=3x+7 ... start with the first equation


x-9=3x+7 ... plug in y = x-9


x-9-7=3x


-9-7=3x-x


-16=2x


2x=-16


x=-16/2


x=-8


------------------------------


Now use this to find y


y=x-9


y=-8-9


y=-17

So the solutions are x=-8 and y=-17